Find the slope of each line segment and interpret the results?
(62,7000) (64,7000) (67,10000) (70,12400)


Answer by Alan3354(69443)

Find the slope of each line segment and interpret the results?
(62,7000) (64,7000) (67,10000) (70,12400)
--------------
Label the points A, B, C, D in order
AB's slope = 0
----------------
BC: m = (10000 - 70000)/3 = 1000
-------------------
CD: m = (12400 - 10000)/3 = 800
---------
DA: m = (7000 - 12400)/8 = -675